OpenStudy (anonymous):

Find the length of a 40 degree arc in a circle with a radius of 4.

OpenStudy (anonymous):

\[len = 2 \Pi r * (40/360)\]

OpenStudy (shayaan_mustafa):

use formula of arc length. \[\large s=r \theta\]

OpenStudy (shayaan_mustafa):

@fractal5 2*pi*r is the circumference of the circle. I think it is just for arc length.

OpenStudy (anonymous):

Can you explain more please?

OpenStudy (anonymous):

@Shayaan_Mustafa multiplying the circumference by the part of the arc, that is 40 deg should give the length of that arc

OpenStudy (shayaan_mustafa):

oh yes. you did so. sorry i just saw 2*pi

OpenStudy (anonymous):

@Brooke_Tessa_Savoie if you know the circumference of the circle, which is like an arc of 360 deg, you can find the length of arc given the angle. Hope that makes sense

OpenStudy (shayaan_mustafa):

@Brooke_Tessa_Savoie http://www.mathwarehouse.com/trigonometry/radians/s=r-theta-formula-equation.php

OpenStudy (shayaan_mustafa):

remember in s=r*theta. theta is in radians.

OpenStudy (anonymous):

I dont know the circumference..

OpenStudy (shayaan_mustafa):

actually you are given with a part of the circumference. and that small part is known as arc.

OpenStudy (anonymous):

You have to find the circumference from the radius C=2Πr

OpenStudy (anonymous):

okay thanks. so the circum=25.12 so 25.12 ( 40/360) ?

OpenStudy (anonymous):

so Now how do I finish solving it?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

just do the math answer = 25.12 * (40 / 360 )

OpenStudy (anonymous):

And how did you get 360?

OpenStudy (anonymous):

because the angle inside a circle is 360 deg, and you just want to find the length of the arc of 40 deg, you basically find the ratio of the arc angle to the whole circle, and multiply by circumference

OpenStudy (anonymous):

It's like cross multiplication If 360 deg gives you 25.12 how much will 40 give you

OpenStudy (anonymous):

Okay Im lost and confused.

OpenStudy (anonymous):

40º is 0.7 rad aproximatly so arc length is 0,7*4 = 2.8

OpenStudy (anonymous):

How did you figure out that 40 is 0.7 rad?

OpenStudy (anonymous):

you can mesure angle in degrees or in radians. (Radian is a mesure system where radius of a circle is taken like unit length) It is known that 360º = 2Pi radians So if you have fraction of 360º (40º/360º) it should be same for any type of mesuring system you choose, since you mesure same angle so 40º/360º = x/2Pi from here x=2Pi( 40º/360º) is a angle of 40º given in radians

OpenStudy (anonymous):

Okay well how do you solve it?

OpenStudy (anonymous):

after you have you angle in radians, just multiply it by radius length.

OpenStudy (anonymous):

The answer is 2.79